### v3

# Shaping chaos in bilayer graphene cavities

Abstract
Bilayer graphene cavities where electrons are confined within finite graphene flakes provide an alluring platform not only for the future nanoelectronic devices owing to the tunable energy gap but also for investigating the quantum nature of chaos due to the trigonal warping of their Fermi surface. Here we demonstrate that rotating the cavity boundary relative to the underlying lattice structure drives a quantum transition from nearly integrable dynamics to chaotic regime, observed as a concomitant crossover of eigenvalue statistics and eigenstate profiles. Complementing the full quantum treatment, we examine the classical backbone of this onset of chaos by employing semiclassical ray dynamics. Our results position bilayer graphene cavities as a promising venue for investigating and engineering quantum-chaotic behavior in graphene-based devices.
preprint: APS/123-QED
I Introduction
Most physical systems are nonintegrable and can exhibit chaotic behavior in their classical limit [ott2002chaos]. The huge qualitative differences in integrable and chaotic behavior can be pivotal to the function of quantum devices. Billiard models have long served as canonical model systems for studying such behavior, in which dynamical complexity arises solely from the geometry of the confining boundary [chernov2006chaotic]. In the quantum regime, billiards correspond to idealized cavities governed by the Schrödinger equation with boundary conditions encoding information on the underlying classical motion, thereby providing a direct window into the quantum signatures of chaos [Stockmann_book, Heller_book, Gutzwiller_book, Haake_book, Tabor_book, Casati_book, Nakamura_book]. This paradigm have played an essential role to establish the semiclassical link between chaotic classical dynamics and quantum features, leading the way towards the random matrix theory [mehta2004random] as a general organizing principle for chaotic quantum systems, further culminating in the Bohigas–Giannoni–Schmit [bohigas1984characterization] and Berry [berry1977regular] conjectures. Consequently, billiard systems have become a foundational testbed for various quantum-chaotic phenomena [Stockmann_book, Heller_book, Gutzwiller_book, Haake_book, Tabor_book, Casati_book, Nakamura_book], ranging from the universality of level statistics and their periodic-orbit corrections to wavefunction ergodicity and its notable exceptions, such as quantum scarring.
Nearly all of the aforementioned results are rooted in the Schrödinger dynamics with an isotropic quadratic dispersion, as commonly realized in semiconductor quantum wells acting as soft-wall billiards [Nakamura_book]. However, not all quantum devices obey the Schrödinger equation; some of them obey the Dirac equation. A distinct class of billiards arises for relativistic wave equations with linear dispersion, first exemplified by the neutrino billiard of Berry and Mondragon [berry1987neutrino]. On the other hand, low-energy quasiparticles in a monolayer graphene (MLG) cavity emulate the physics of the Dirac equation [ponomarenko2008chaotic, condado2025exactly, huang2018relativistic], in which pseudospin, valley degrees of freedom, and relativistic boundary conditions fundamentally modify the correspondence between classical dynamics and quantum properties. As a result of being more accessible than a hypothetical neutrino billiard, an extensive amount of investigations has focused on MLG devices to examine the characteristics of a pseudo-relativistic quantum system, such as Klein tunneling [Beenakker_rev.mod.phys_80_1337_2008], the creation and probing of whispering gallery mode resonators [zhao2015creating], spectral statistics and wavefunction properties [geim2007rise, castro2009electronic, PhysRevB.77.085423] and recently the direct visualization of relativistic quantum scars [ge2024direct].
Beyond the extensive interest in MLG [Castro_rev.mod.phys_81_109_2009, DasSarma_rev.mod.phys_83_407_2011, Basov_rev.mod.phys_86_959_2014], bilayer graphene (BLG) has emerged as a highly versatile platform in its own right [mccann2013electronic]. Recent advances have demonstrated exceptional material quality and electrostatic control [Tong_nano.lett_21_1068_2021, Banszerus_nano.lett_20_7709_2020], together with a wealth of novel physical phenomena [Banszerus_nat.commun_13_3637_2022, Garreis_phys.rev.lett_126_147703_2021, Garreis_nat.phys_30_428_2024, Garreis_phys.rev.research_5_013042_2023, Kurzmann_nat.commun_12_6004_2021, Kurzmann_phys.rev.lett_123_026803_2019, Tong_phys.rev.lett_133_017001_2024]. For example, an out-of-plane electric field induces a tunable band gap by breaking layer symmetry, enabling electrostatically defined confinement [mccann2013electronic, McCann_phys.rev.lett_96_086805_2006, Min_phys.rev.b_75_155115_2007, Castro_phys.rev.lett_99_216802_2007, Zhang_nature_459_820_2009]. Alternatively, the necessary confinement can be created through lithographic methods [see, e.g., Refs. [Bell_nanotechnology_20_455301_2009, Cong_j.phys.chem.c_113_6529_2009, rubio_phys.chem.chem.phys_19_8061_2017]], where the BLG is either cut or grown into a flake of desired geometry, producing hard-wall confinement reminiscent of a quantum billiard. In contrast to MLG, BLG lacks relativistic Dirac dynamics at low energies but instead hosts distinct phenomena, including Zitterbewegung-like motion arising from interlayer coupling [Bao_phys.rev.e_92_012918_2015]. Even though the low-energy dispersion is quadratic, it is strongly modified by trigonal warping [mccann2013electronic].
The pronounced role of trigonal warping in BLG has been demonstrated under a wide range of conditions [Garreis_phys.rev.lett_126_147703_2021, Gold_phys.rev.lett_127_046801_2021, Kechedzhi_phys.rev.lett_98_17680_2007, Kaladzhyan_phys.rev.b_104_235425_2021, Garreis_nat.phys_30_428_2024], with scanning tunneling microscopy providing quasi-direct access to its anisotropic Fermi surface. This anisotropy renders the electronic dynamics highly sensitive to electrostatic gating [ohta2006controlling, oostinga2008gate] and strain [PhysRevB.84.041404], as well as to lattice orientation. Notably, gate tuning alone can induce a coexistence of regular and chaotic dynamics even in highly symmetric BLG quantum dots [seemann2023gate]. Compared to MLG and common two-dimensional electron gas structures, trigonal warping allows the semiclassical dynamics to be tuned with far greater flexibility. In fact, semiclassical ray-dynamics and Poincaré analyses have revealed phase-space structures and periodic-orbit stability properties that are qualitatively distinct from those of conventional billiards [seemann2023gate, seemann2025complex, schoeppl2025anisotropic]. Furthermore, the same anisotropy underlies electron-optical effects such as Veselago lensing, underscoring the unusual transport properties of BLG cavities [cheianov2007focusing].
While the features of of BLG have motivated investigation in transport and device physics [PhysRevB.81.195406, san2014stacking], the implications remain largely uncharted in the context of quantum chaos, furthermore almost entirely confined to semiclassical perspective. The quantum mechanical properties of BLG cavities, such as their eigenstate morphology, spectral statistics, and the extent to which the semiclassical features persist or break down at the quantum level, are essentially unknown. Addressing these issues is crucial for understanding BLG cavities not only as electron-optical devices [cheianov2007focusing] but also as a platform for systematically investigating quantum-chaotic phenomena in strongly anisotropic materials.
Here, we demonstrate that boundary–lattice misalignment within BLG cavities induces a transition from the near-integrability to chaos, reflected in both eigenvalue statistics and eigenstate anatomy, while semiclassical ray dynamics exposes the classical phase-space restructuring responsible for this behavior that aligns with the prior work by Hentschel and co-workers [hentschel2025ray, seemann2025complex]. In particular, we investigate the interplay of the discrete nature of the lattice and the cavity geometry, going beyond widely used continuum models, such as utilized in Ref. [schoeppl2025anisotropic]. Taken all together, we present a step toward the systematic investigation and design of chaos in BLG–based devices.
Figure 1: Illustration of the tight-binding model, Fermi surface, and cavity geometry. (a) Schematic diagram of the BLG tight-binding model showing atoms and hopping. (b) K valley of BLG’s band structure, where the white curve represents the Fermi surface at $E=0.2\,\mathrm{eV}$, showing trigonal warping due to interlayer hopping $\gamma_{3}$. (c) and (d) show diagrams of BLG cut into a hexagonal shape, where purple indicates the dimer atoms $B_{1}$ and $A_{2}$, i.e., the vertically aligned sites that are directly coupled between the two layers, and red and blue indicate the non-dimer atoms $A_{1}$ and $B_{2}$. In (c), the boundary is aligned with the inner lattice, and the point group of the atoms is $D_{3d}$. In (d), the boundary is rotated by $15^{\circ}$ relative to the inner lattice, and the point group of the atoms is $S_{6}$. The cavities we study contain millions of atoms.
II Model
We investigate an AB-stacked BLG cavity with regular hexagonal geometry, in order to elucidate the influence of trigonal warping and lattice–boundary misalignment. More specifically, the electronic structure of this cavity is described by the standard Slonczewski–Weiss–McClure tight-binding Hamiltonian [mccann2013electronic]:
$\mathcal{H}$ $=\sum_{\langle i,j\rangle_{\mathrm{intra}}}\gamma_{0}\,c_{i}^{\dagger}c_{j}+\sum_{\langle B_{1},A_{2}\rangle}\gamma_{1}\,c_{B_{1}}^{\dagger}c_{A_{2}}+\sum_{\langle A_{1},B_{2}\rangle}\gamma_{3}\,c_{A_{1}}^{\dagger}c_{B_{2}}$
$+\sum_{\langle A_{1},A_{2}\rangle}\gamma_{4}\,c_{A_{1}}^{\dagger}c_{A_{2}}+\sum_{\langle B_{1},B_{2}\rangle}\gamma_{4}\,c_{B_{1}}^{\dagger}c_{B_{2}}+\text{H.c.},$ (1)
which includes the dominant intralayer hopping $\gamma_{0}$, the interlayer dimer coupling $\gamma_{1}$, and the skew interlayer hoppings $\gamma_{3}$ and skew interlayer hopping between same-sublattice sites $\gamma_{4}$. In the Hamiltonian above, $c_{i}^{\dagger}$ ($c_{i}$) creates (annihilates) an electron on site $i$, and $\langle\cdot,\cdot\rangle$ denotes the corresponding in-plane or interlayer nearest-neighbor pairs as indicated in Fig. 1(a). Furthermore, $A_{1}$ and $B_{1}$ in Fig. 1(a) refer to the two sublattice atoms in the lower graphene layer, while $A_{2}$ and $B_{2}$ refer to the two sublattice atoms in the upper layer. Non-dimer sites in the top and bottom layers are shown in blue and red, respectively, while vertically aligned interlayer dimer sites are marked in purple.
The hopping parameters are chosen in accordance with established literature [mccann2013electronic] (see Materials and Methods for details). Whereas dominant in-plane hopping $\gamma_{0}$ defines the Dirac-like dispersion within each layer and the vertical interlayer coupling $\gamma_{1}$ hybridizes the dimer sites ($B_{1}$$A_{2}$) yielding the characteristic low-energy parabolic bands, we are most interested in the skew parameter $\gamma_{3}$ which introduces trigonal warping and hence breaks the isotropy of the energy spectrum near the neutrality point. Fig. 1(b) shows a trigonally warped Fermi surface of BLG stemming from the non-dimer coupling parametrized by $\gamma_{3}$. Finally, same-sublattice interlayer hopping $\gamma_{4}$ couples $A_{1}$$A_{2}$ and $B_{1}$$B_{2}$ sites, lifting up electron–hole symmetry and only slightly modifying the band curvature.
We focus on the energy window $\Delta E=0.18$$0.25\,\mathrm{eV}$, where the Fermi surface exhibits pronounced trigonal warping. As displayed in Fig. 1(b), the $K$-valley Fermi surface becomes strongly anisotropic at energies a few hundred meV above the quadratic band-touching point. Within this energy window, the typical wavelength of a quantum state, $\lambda\approx 20\,\mathrm{nm}$, remains much larger than the lattice constant $a\approx 0.246\,\mathrm{nm}$. Moreover, the characteristic system size $L\approx 800\,\mathrm{nm}\gg\lambda$ places the studied system firmly in the semiclassical regime, justifying the semiclassical analysis employed below. Crucially, we have verified that both the numerical results and the corresponding conclusions below are qualitatively robust against variations in system size (the analysis is carried out in Supplementary Material).
To directly probe the interplay between the underlying and cavity boundary, we rotate the original hexagonal BLG cavity shown in Fig. 1(c) by an angle $\theta$, as demonstrated in Fig. 1(d). However, the sixfold rotational symmetry of the lattice renders the atomic configuration $60^{\circ}$-periodic in the rotation angle $\theta$. Furthermore, since the configurations at the angles $\theta$ and $60^{\circ}-\theta$ are equivalent, it is hence sufficient to restrict our consideration to the interval $0^{\circ}\leq\theta\leq 30^{\circ}$. The limiting cases $0^{\circ}$ and $30^{\circ}$ structurally correspond to boundary-aligned zigzag and armchair terminations, respectively; whereas intermediate angles produce misaligned edges with mixed zigzag–armchair character.
III Results
We compute eigenvalues and eigenstates of the Hamiltonian in (1) for the distinct angles $0^{\circ}\leq\theta\leq 30^{\circ}$, and then quantify the degree of chaos within the energy window $\Delta E$. The simulation details are provided in Materials and Methods and Supplementary Material. However, before delving into these results, we briefly highlight two underlying key aspects about the chosen system.
First, we address the symmetry of the system (see Materials and Methods for details). In the case of the unrotated cavity ($\theta=0$) shown in Fig. 1(c), the boundary is aligned with the crystalline axes, and the Hamiltonian respects the $D_{3d}$ point-group symmetry, which further decomposes into the six irreducible representations $A_{1g}$, $A_{2g}$, $A_{1u}$, $A_{2u}$, $E_{g}$, and $E_{u}$. By contrast, for a representative rotated configuration $(\theta=15^{\circ})$ presented in Fig. 1(d)), the boundary orientation breaks the orginal mirror symmetries, downgrading the point group to $S_{6}$, with four sectors $A_{g}$, $A_{u}$, $E_{g}$, and $E_{u}$.
Second, we note that a classical hexagonal billiard is an archetypal example of a pseudointegrable system [richens1981pseudointegrable, lozej2024intermediate]: its interior angles are rational multiples of $\pi$, so trajectories evolve on higher-genus phase-space surfaces rather than on invariant tori as in a fully integrable system. On the quantum side, such pseudointegrable billiards display hybrid spectral statistics, interpolating between regular and chaotic behavior, but are frequently characterized by the semi-Poisson distribution [berry1984semiclassical, PhysRevLett.62.2769, bogomolny1999models, giraud2004intermediate, lozej2024intermediate, biswas1990quantum] (see also Materials and Methods). Our BLG cavity, however, differs from the idealized classical and quantum billiards due to the trigonal warping, and the discrete lattice which removes the sharp boundary corners causing classical pseudointegrable dynamics. Therefore, it is not a priori clear whether pseudointegrability carries over to our system. As we show next, the BLG cavity does inherit this feature to some degree, and we also explore its manifestation in the eigenfunction profiles that is a much less studied aspect compared to spectral statistics.
III.1 Level Statistics
We begin by estimating the degree of chaos in terms of the nearest-neighbor level spacing distribution and the spectral rigidity $\Delta_{3}(L)$ which are computed within each symmetry-resolved sector, probing the short- and long-range spectral correlations, respectively. Additional quantification is provided by the average gap ratio $\langle\tilde{r}\rangle$ and the average spectral rigidity $\langle\Delta_{3}\rangle$. Figure. 2 summarizes these measures for the unrotated cavity and for a cavity rotated by $\theta=15^{\circ}$. We further classify the (sub)spectra following the spirit of the Bohigas–Giannoni–Schmit conjecture [bohigas1984characterization]: a regular spectrum is associated with Poisson statistics, whereas a chaotic spectrum conform to the Gaussian Orthogonal Ensemble (GOE) or Gaussian Unitary Ensemble (GUE), depending on the presence or absence of time-reversal symmetry, respectively. The latter is also referred as Wigner-Dyson behavior. Our quantum-chaos estimations are detailed in Materials and Methods.
Figure 2: Level Statistics of unrotated cavities and rotated BLG cavities. (a) and (b) show the level spacing statistics and spectral rigidity for the unrotated cavity, with the corresponding $\langle\tilde{r}\rangle$ values and averaged rigidity indicated in each panel. The red dashed line denotes the Poisson distribution, the blue solid line represents the Wigner–Dyson distribution, and the green dash-dotted line shows the fitted semi-Poisson distribution. (c) and (d) present the same analyses for the cavity rotated by $15^{\circ}$. The purple solid line represents the GUE distribution, and the orange dash-dotted line denotes the fitted GOE–GUE distribution.
The unrotated hexagonal cavity exhibits a pronounced sector-to-sector separation: two one-dimensional sectors, $A_{1u}$ and $A_{2g}$, are nearly uncorrelated (Poisson-like), whereas the remaining sectors display intermediate, semi-Poisson statistics characteristic of pseudointegrable polygonal billiards. This sector dependence reflects the different boundary constraints imposed by the symmetry projections, which can lead to distinct effective dynamics within the same device.
The Poisson-like behavior of the $A_{1u}$ and $A_{2g}$ sectors originates from a symmetry reduction of the boundary-value problem. In this sector, the parity constraints enforce nodal lines that act as effective Dirichlet boundaries, so the hexagon reduces to an equilateral-triangle fundamental domain. Since the equilateral triangle is integrable, this sector naturally exhibits near-Poisson statistics, consistent with the standard quantum billiard correspondence. Atomistic zigzag details do not change this large-scale geometry.
[…]
The rotated cavity shows a qualitatively different behavior. As illustrated by Fig. 2(c,d) for $\theta=15^{\circ}$, the $A$ sector remains intermediate with $\langle\tilde{r}\rangle=0.516$, while the $E$ sector shifts toward Wigner–Dyson correlations with $\langle\tilde{r}\rangle=0.558$ and a strongly suppressed exponential tail. The data are well described by a mixed GOE–GUE interpolation [mehta2004random](see Materials and Methods for the fitting form), demonstrating that lattice–boundary misalignment removes the near-integrable channel present at $\theta=0^{\circ}$ and drives the spectrum toward fully chaotic statistics. The rigidity shows a consistent trend: $\langle\Delta_{3}\rangle$ decreases to $0.31$ in $A$ and $0.24$ in $E$, approaching the GOE/GUE limits.
Figure 3: Angle dependence of $\langle\tilde{r}\rangle$ and $\langle\Delta_{3}\rangle$ on $\theta$. (a) and (b) present the dependence of the $r$-value and the average spectral rigidity on the rotation angle $\theta$ for the $A$ and $E$ subspaces, respectively. In the AA subspaces, two lines are shown: for the unrotated cases $\theta=0^{\circ}$ and $30^{\circ}$, the blue solid line corresponds to the pseudointegrable sector, while the purple dashed line represents the integrable sector.
Figure 4: The wavefunctions and their associated statistical properties. (a1)–(a3) representative probability density distributions $|\psi(\mathbf{r})|^{2}$ of eigenstates in unrotated cavities for different symmetry subspaces. (a4) and (a5) the corresponding results for the rotated case. (b1)–(b5) the associated distributions $|\tilde{\psi}(\mathbf{k})|^{2}$ in momentum space with trigonal warped Fermi surface at $K$ and $K^{\prime}$. (c1)–(c5) the statistical behavior of the correlation length $l$ for the unrotated and rotated cavities across the symmetry subspaces.
Fig. 3 summarizes the evolution with rotation angle. In the $A$ sector, one branch remains in the intermediate (semi-Poisson-like) regime across all angles, while the branch that is Poisson-like at $\theta=0^{\circ}$ and $30^{\circ}$ rapidly increases toward Wigner–Dyson values as $\theta$ departs from these commensurate alignments. In the $E$ sector, $\langle\tilde{r}\rangle$ increases more steadily and approaches the GOE/GUE range over a broad interval of angles. The averaged rigidity $\langle\Delta_{3}\rangle$ exhibits the same overall trend, shown in Fig. 3(b), indicating that the enhancement of chaotic correlations is robust and does not require fine tuning of $\theta$.
III.2 Eigenstates
To connect the spectral crossover to wavefunction structure, we study the eigenstates and quantify how lattice–boundary misalignment modifies their spatial correlations. The structure of eigenstates reflects the nature of the underlying classical dynamics. Integrable systems exhibit regular standing-wave patterns, whereas chaotic systems follow Berry’s random wave model and may host quantum scars associated with unstable classical orbits [mehta2004random, berry1977regular, heller1984bound]. Pseudointegrable systems interpolate between these limits and often show symmetry-dependent eigenstate structures [biswas1990quantum].
To quantify how rotation modifies the eigenstates, we characterize their spatial structure using an autocorrelation function, from which we extract a correlation length $l$ as a measure of spatial coherence (see Materials and Methods). The correlation length $l$ measures the typical length scale over which the wavefunction remains correlated: smaller $l$ indicates more chaotic, spatially uncorrelated states, while larger $l$ corresponds to more regular behavior. For a fully chaotic system, Berry’s random wave model predicts $l$ to be equal to the wavelength $\lambda$ of the wavefunction [berry1977regular]. In our system, $\lambda\approx 20\,\mathrm{nm}$ at energies around $0.2\,\mathrm{eV}$. We compute statistical averages of $l$ from 1,056 eigenstates near this energy to assess the degree of chaos as a function of rotation angle.
We present representative eigenstates from each symmetry subspace in Fig. 4. The real-space probability density distributions are shown in Figs. 4(a1)–(a5), while the corresponding momentum-space structures, obtained from the Fourier transform $\tilde{\psi}(\mathbf{k})=\int d^{2}\mathbf{r}\,e^{-i\mathbf{k}\cdot\mathbf{r}}\,\psi(\mathbf{r})$, are displayed in Figs. 4(b1)–(b5), together with the trigonal-warped Fermi surfaces in the $K$ and $K^{\prime}$ valleys.
For the unrotated cavity, the $A_{1u}$ and $A_{2g}$ sectors exhibit regular standing-wave patterns (see Fig. 4(a1)), characteristic of integrable eigenstates [berry1981regularity]. Correspondingly, its momentum-space density is localized at discrete points on the Fermi surface (see Fig. 4(b1)), reflecting the dominance of a small number of plane-wave components. In contrast, eigenstates from the $A_{1g}$ and $A_{2u}$, $E_{u}$ and $E_{g}$ sectors display more irregular spatial structures (see Figs. 4(a2),(a3)), accompanied by broader momentum-space distributions involving a larger set of wave vectors (see Figs. 4(b2),(b3)), consistent with their pseudointegrable nature.
Representative eigenstates from the rotated cavities, shown in Figs. 4(a4),(a5), exhibit highly irregular real-space density profiles without clear signatures of periodic orbits. Their momentum-space densities (see Figs. 4(b4),(b5)) form nearly continuous distributions along the trigonal-warped Fermi surface, indicating random-wave–like behavior. Unlike Berry’s random wave model with isotropic wave vectors, the randomness here arises from plane waves constrained to the warped Fermi surface.
The statistical analysis of the correlation length is summarized in Figs. 4(c1)–(c3). In the unrotated cavity, all symmetry subspaces yield average correlation lengths much larger than the wavelength $\lambda=20\,\mathrm{nm}$, indicating the absence of full chaos. The integrable $A_{1u}$ and $A_{2g}$ sectors show the largest value, $\langle l\rangle=531\,\mathrm{nm}$, while the pseudointegrable $A_{1g}$ and $A_{2u}$, $E_{u}$ and $E_{g}$ sectors exhibit smaller values, $\langle l\rangle=299\,\mathrm{nm}$ and $169\,\mathrm{nm}$, respectively, yet still significantly exceeding $\lambda$.
The change in the correlation length statistics upon rotation is striking. We consider a cavity rotated by $\theta=15^{\circ}$, as shown in Figs. 4(c4) and (c5). For both the $A$ and $E$ sectors, the average correlation length decreases dramatically compared to the unrotated case, reaching $\langle l\rangle_{A}=39.2\,\mathrm{nm}$ and $\langle l\rangle_{E}=23.3\,\mathrm{nm}$. Notably, in the $E$ subspace, the correlation lengths of most eigenstates cluster around the wavelength $\lambda=20\,\mathrm{nm}$, indicating the emergence of spatially uncorrelated, random-wave–like eigenstates.
The statistical behavior of the correlation length for a wider range of rotation angles is presented in the Supplementary Material. In addition, we analyze the momentum space distribution in Supplementary Material, which shows that rotation leads the wavefunctions to be composed of a larger number of plane waves with different wavevectors. These results indicate that rotation of the BLG cavity induces quantum chaos by significantly suppressing spatial correlations and increasing the number of contributing plane wave components for the eigenstates.
IV Semiclassical Dynamics
To provide an intuitive dynamical picture of how trigonal warping reshapes the cavity behavior, we complement the tight-binding analysis with a simplified semiclassical ray model. Here, we isolate the geometric effects of the warped dispersion and the resulting anisotropic group velocity. We therefore adopt a continuum bulk description based on the standard four-band Hamiltonian for AB-stacked bilayer graphene [mccann2013electronic], used here as a minimal description of warping-induced anisotropy in the bulk dispersion. This model intentionally omits atomistic edge details and serves as a heuristic tool for visualizing how lattice–boundary misalignment modifies the effective cavity dynamics.
In this framework, the cavity is modeled as an anisotropic Minkowski billiard [gutkin2002billiards]: rays propagate with the band velocity $\mathbf{v}(\mathbf{k})=\nabla_{\mathbf{k}}E(\mathbf{k})$ along a warped constant-energy contour, and each boundary collision is implemented by enforcing conservation of energy and the tangential momentum component (see Supplementary Material for the technical details).
Unlike geometric mirror reflection, this rule is momentum dependent and highly anisotropic. Nevertheless, because the warped Fermi contour remains strictly convex in our energy window, the resulting reflection map is smooth and area preserving. For polygonal billiards, the Lyapunov exponent is always zero. Because the boundary has no curvature and therefore cannot generate exponential separation of nearby trajectories. While certain polygons may be ergodic, they do not exhibit classical chaos. In our BLG ray-dynamics, trigonal warping introduces a momentum-dependent and highly anisotropic reflection rule. Although the angular magnification in Fig 5 (d) displays pronounced fluctuations and sharp peaks for specific incidence angles, our numerical evaluation still yields a maximal Lyapunov exponent $\lambda_{\max}=0$. Thus, trigonal warping does not generate exponential sensitivity in the semiclassical dynamics. However, $\lambda_{\max}=0$ does not mean that the dynamics is unchanged. As shown below, the anisotropic reflection map can strongly modify the ergodicity of the dynamics.
Figure 5: Poincaré sections constructed by recording, for a single trajectory crossing a chosen reference edge, the boundary coordinate $s$ (position along the edge) and the incidence angle $\theta$ (relative to the boundary normal). (a) With the boundary aligned to crystalline $C_{3}$ axis, the map collapses onto a few invariant curves characteristic of pseudo-integrable motion. (b) After rotating the cavity by $15^{\circ}$, the map spreads quasi-ergodically over the accessible phase space.(c) For comparison, a mirror reflecting irrational hexagon shows uniformly filled Poincaré points, demonstrating true ergodicity without crystalline commensurability. (d) The angular magnification $|d\phi_{\rm out}/d\phi_{\rm in}|$ for BLG’s anisotropic reflection (blue) compared with the constant unit magnification of specular mirror reflection (red dashed). The variations in BLG magnification quantify the underlying warping responsible for the structures in Figs. 5 (a) and (b).
The Poincaré sections in Fig. 5 are constructed by recording the arclength coordinate $s$ and incidence angle $\theta$ whenever a ray intersects a fixed reference edge of the hexagon, over $5\times 10^{5}$ collisions. When the boundary is aligned with a crystalline $C_{3}$ axis, the warped reflection rule is commensurate with the boundary symmetry. As a result, the warped reflection rule effectively reduces to ordinary specular reflection [seemann2023gate]. Successive reflections recur along a small set of preferred directions. Consecutive reflections therefore repeat only a small set of incidence angles, and the resulting Poincaré section collapses into a few discrete lines rather than filling the available phase space (see Fig. 5(a)). The motion is confined to these invariant lines, consistent with pseudo-integrable dynamics.
After rotating the cavity by $15^{\circ}$, this commensurability is lost. The warped reflection rule becomes incommensurate with the boundary, and incidence angles no longer recur coherently. The resulting Poincaré section displays a dense but structured phase-space pattern consistent with the quasi-ergodic regime discussed by Wang et al [wang2022statistical]: the trajectory becomes dense in phase space, but the point cloud remains pronounced non-uniform even after long evolution times. Compared to Fig. 5(c), bands of enhanced density and filamentary streaks persist across the section, and the occupancy of phase space shows clear variations in the coordinates $(s,p)$ (see Fig. 5(b)). This slow, non-uniform phase-space filling provides a classical counterpart to the spectral crossover discussed above.
For comparison, Fig. 5(c) shows a mirror-reflecting irrational hexagon. The Poincaré points fill the allowed domain uniformly, demonstrating genuinely ergodic behavior. Compared to Fig. 5(b), the trajectory fills the phase space uniformly after only a short time.
V Discussion and Conclusion
The emergence of quantum chaos upon misaligning the boundary and the lattice can be understood from a semiclassical perspective. We emphasize that the chaos here refers to quantum chaos, indicated by Wigner–Dyson spectral statistics and random wave eigenstates. Although we use a simplified semiclassical ray-dynamics construction, this model is intended to highlight the geometric consequences of the warped dispersion. No simple semiclassical description fully captures the atomistic boundary physics of the lattice model.
The simplified semiclassical picture nonetheless provides a meaningful interpretation of the observed transition. The rotated cavity exhibits quasi-ergodic phase space spreading, and such semiclassical ergodicity is known to be sufficient to generate Wigner–Dyson statistics even in the absence of strong classical chaos or positive Lyapunov exponents. Indeed, studies of non-hyperbolic systems, such as the quantum triangle map [wang2022statistical], triangular cavities [lozej2022quantum], and parabolic triangle maps [degli2005semi], have shown that classical quasi-ergodicity alone is sufficient to produce GOE or GUE level statistics in the semiclassical limit, highlighting the role of classical mixing and ergodic properties in shaping quantum spectral behavior.
In addition, for a rotated BLG cavity defined on a lattice, one must distinguish geometric incommensurability from atomistic edge roughness. When the cavity is rotated relative to the underlying lattice, the boundary becomes a mixture of zigzag and armchair segments and develops kinks. Since mixed edges are known in MLG to modify intervalley scattering and consequently affect spectral correlations and eigenstate properties in specific cavity geometries [yu2016gaussian, hagymasi2017interaction], one may worry that edge effects could play a dominant role in BLG. To assess the role of edge roughness, we perform two sets of control tests. First, we introduce artificial kinks on an unrotated hexagon and find that increasing the number of kinks leads to a slow increase of $\langle\tilde{r}\rangle$. Second, we rotate monolayer cavity, where trigonal warping is absent at low energies and the symmetry is reduced from $D_{6}$ to $C_{6}$ upon rotation. In MLG, $\langle\tilde{r}\rangle$ increases slightly, but the spectrum remains semi-Poisson even up to $10^{\circ}$ (see Supplementary Information). This comparison indicates that edge roughness provides a gradual, angle-dependent contribution to the spectral correlations.
In BLG, by contrast, trigonal warping provides a much stronger mechanism that can drive a rapid transition. A rotation as small as $0.1^{\circ}$, corresponding to shifting the lattice by only one kink and producing negligible geometric deformation, is already sufficient to drive the spectrum to Wigner–Dyson statistics. At larger rotation angles, edge effects become progressively more important, since the increased armchair content enhances intervalley scattering and the semiclassical reflection rule is no longer expected to apply [PhysRevB.84.075468]. We therefore view the onset of quantum chaos in rotated BLG cavities as the combined outcome of (i) a dominant trigonal-warping mechanism that rapidly destroys the commensurability between the boundary and the bulk dynamics, and (ii) additional edge roughness effects that become increasingly relevant at larger rotation angles.
An important exception is the $30^{\circ}$, where the lattice and boundary become commensurate and edge physics becomes dominant. In this geometry, the spectrum shifts back toward quasi-integrable behavior. The boundary is closer to a uniform armchair termination, where intervalley physics is enhanced. Because our semiclassical model neglects atomistic edge conditions, its bulk-mismatch criterion is not expected to capture this edge-dominated symmetry restoration.
The BLG cavity model in our study can be experimentally realized. A hexagonal graphene cavity can be experimentally realized via CVD growth, and the rotated cavity model considered in our work corresponds to the hexagon with kinks [2013PNAS_CVD]. In addition, such cavities can be obtained either by patterning BLG into discrete shapes, creating hard-wall confinement [ponomarenko2008chaotic, barreiro2012quantum, hagymasi2017interaction], or by using electrostatic gates to define a soft confinement potential, allowing tunable control over the dot occupancy and tunnel barriers [subramaniam2012wave, goossens2012gate, kurzmann2019charge]. Low temperature transport measurements, such as Coulomb blockade and excited state spectroscopy, can probe energy level statistics and wavefunction distributions, enabling experimental access to signatures of quantum chaotic behavior [banszerus2020single]. These approaches offer complementary ways to explore and control quantum chaos in BLG-based systems.
For future studies of BLG cavities, one can focus on controlling quantum behavior by tuning various parameters of the system. One promising direction is to introduce an external electric field, which can modify electron dynamics and selectively generate or tune specific wavefunction patterns. The interplay among the electric field, cavity geometry, and trigonal warping may give rise to a rich variety of dynamical regimes and quantum phenomena. In particular, it may enable the controlled manipulation of superscars in classically pseudointegrable cavities [2004PRLBogomolny, 2006PRL_Superscar_Bogomolny]. Other factors, such as an applied magnetic field or an interlayer potential difference, can also be used to tune the system by altering its symmetry and internal electronic dynamics.
In conclusion, we have shown that misalignment between the BLG lattice and the hexagonal boundary fundamentally reshapes both the spectral and eigenstate properties of the cavity. Rotating the boundary removes the high symmetry subspaces in the unrotated cavity, produces Wigner–Dyson distribution of level spacing, and drives the eigenstates’ correlation lengths toward the wavelength scale, signaling the emergence of fully chaotic eigenstates. The semiclassical analysis confirms that this transition originates from the loss of commensurability between the warped Fermi surface and the polygonal boundary, which generates ergodicity in phase space and correspondingly causes quantum chaos. Overall, our results demonstrate that the boundary-lattice interplay provides an effective and experimentally accessible route for engineering and controlling quantum chaos in BLG cavities, with direct relevance to controllable electron transport in graphene and other mesoscopic systems [Velasco_2012, RevModPhys2015_cao, PhysRevEconductance_2015].
VI Materials and Methods
VI.1 Numerical implementation of eigenvalue and eigenstate simulations
All numerical simulations presented in this work are performed using the pybinding Python package, which provides a flexible and efficient framework for tight-binding modeling of two-dimensional materials [moldovan2020pybinding]. Working with an atomistic tight-binding Hamiltonian on the physical honeycomb bilayer lattice avoids discretizing a continuum model on a Cartesian real-space grid, which can introduce artificial square-lattice ($D_{4}$) anisotropies that contaminate symmetry-sensitive spectral statistics [schoeppl2025anisotropic]. For the same reason, our calculations do not suffer from spurious fermion-doubling artifacts associated with finite-difference discretizations of continuum Dirac theories [nielsen1981absence].
VI.2 Tight-binding model parameters
For the tight-binding model of BLG, we adopt the commonly used parameters for AB-stacked layers. The lattice constant for BLG is $a=0.246~\mathrm{nm}$, and the interlayer distance is $d=0.335~\mathrm{nm}$. The intralayer nearest-neighbor hopping is $t_{0}=3.16$ eV, the interlayer vertical hopping between dimer sites is $t_{1}=0.381$ eV, and the skew interlayer hoppings are $t_{3}=0.38$ eV and $t_{4}=0.14$ eV [PhysRevB.81.195406].
The circumradius of the cavity is set to $r=400\,\textrm{nm}$ that corresponds to a characteristic size $L\approx 800\,\textrm{nm}$. In our numerical calculations, the hexagonal cavities contain on the order of $10^{6}$ lattice sites, but we are able to to resolve fully quantum spectrum in this regime where semi-classical or continuum approximations would commonly be applied.
VI.3 Symmetry Analysis
The symmetry of the tight-binding Hamiltonian differs between the two configurations we study. When the boundary is aligned with the crystalline axes, the tight-binding Hamiltonian respects the full $D_{3d}$ symmetry of the BLG lattice. This group contains a threefold rotation $C_{3}$, inversion $\mathcal{I}$, and three vertical mirror operations $\sigma_{v}$.
Since $\mathcal{H},\mathcal{R}_{2\pi/3}]=0$, the eigenstates of $\mathcal{H}$ can be chosen as simultaneous eigenstates of the $120^{\circ}$ rotation operator $\mathcal{R}_{2\pi/3}$, namely
$\mathcal{R}_{2\pi/3}\psi_{n}^{(m)}(\mathbf{r})=e^{i2\pi m/3}\psi_{n}^{(m)}(\mathbf{r}),\quad m=0,\pm 1.$ (2)
The rotation-invariant sector $m=0$ further splits under inversion $\mathcal{I}$ and mirror reflection $\sigma_{v}$ into four one-dimensional irreducible representations. To make these parity properties explicit, we label the corresponding eigenstates as $\psi_{n}^{(0,p_{i},p_{\sigma})}$, where $p_{i}=\pm 1$ and $p_{\sigma}=\pm 1$ denote the eigenvalues under inversion and mirror operations, respectively:
$\mathcal{I}\,\psi_{n}^{(0,p_{i},p_{\sigma})}=p_{i}\,\psi_{n}^{(0,p_{i},p_{\sigma})},\qquad\sigma_{v}\,\psi_{n}^{(0,p_{i},p_{\sigma})}=p_{\sigma}\,\psi_{n}^{(0,p_{i},p_{\sigma})}$ (3)
In this notation, $(p_{i},p_{\sigma})=(+,+)$, (+,)(+,-), (,+)(-,+), and (,)(-,-) correspond to the irreducible representations $A_{1g}$, $A_{2g}$, $A_{1u}$, and $A_{2u}$, respectively. The correspondence between the eigenvalues of the symmetry operations ($m$, $p_{i}$, and $p_{\sigma}$) and the irreducible representations is summarized in Table 1 (left).
Table 1: Symmetry sectors used for $D_{3d}$ (left) and $S_{6}$ (right). Listed are the irreducible subspace’s dimension, rotation sector $m$, and inversion/mirror parities $p_{i}/p_{\sigma}$.
Group Irrep Dim. $m$ $p_{i}$ $p_{\sigma}$
$D_{3d}$ $A_{1g}$ 11 0 +1+1 +1+1
$A_{2g}$ 11 0 +1+1 1-1
$A_{1u}$ 11 0 1-1 +1+1
$A_{2u}$ 11 0 1-1 1-1
$E_{g}$ 22 $\pm 1$ +1+1 0
$E_{u}$ 22 $\pm 1$ 1-1 0
Group Irrep Dim. $m$ $p_{i}$
$S_{6}$ $A_{g}$ 11 0 +1+1
$A_{u}$ 11 0 1-1
$E_{g}$ 22 $\pm 1$ +1+1
$E_{u}$ 22 $\pm 1$ 1-1
When the boundary is rotated by an angle $\theta$, mirror symmetry is broken, and the symmetry group of the tight-binding Hamiltonian reduces to $S_{6}$. The corresponding symmetry properties and irreducible representations are listed in Table 1 (right). Based on the symmetry analysis, we can study the eigenstates independently in each subspace.
VI.4 Level Statistics
To characterize the spectral properties for each orientation angle $\theta$, we compute the ordered eigenvalues $\{E_{i}\}$ of the tight-binding Hamiltonian and analyze their fluctuations after unfolding. Nearest-neighbor spacings are defined as $s_{i}=E_{i+1}-E_{i}$, and the resulting spacing distribution $P(s)$ is compared with the standard Poisson and Wigner–Dyson forms of random-matrix theory [bohigas1984characterization]. We evaluate the ratio of consecutive spacings, $r_{i}=\frac{s_{i+1}}{s_{i}}$, $\tilde{r}_{i}=\min(r_{i},1/r_{i})$, which takes the characteristic values $\langle\tilde{r}\rangle_{\mathrm{P}}\simeq 0.386,\langle\tilde{r}\rangle_{\mathrm{GOE}}\simeq 0.531,$ and $\langle\tilde{r}\rangle_{\mathrm{GUE}}\simeq 0.600$ [atas2013distribution].
Long-range correlations of spectrum are quantified using the spectral rigidity [dyson1963statistical], which is defined as
$\Delta_{3}(L)=\frac{1}{L}\min_{\{A,B\}}\int_{E}^{E+L}\!\!\big[N(E^{\prime})-AE^{\prime}-B\big]^{2}\,dE^{\prime}$ (4)
where $N(E)$ is the unfolded staircase function. While integrable spectra show a linear increase of $\Delta_{3}(L)$ and chaotic spectra exhibit the logarithmic behavior predicted by random-matrix theory, the orientation-dependent evolution of $\Delta_{3}$ provides a complementary measure of long-range correlations in BLG cavities. To quantify the degree of chaos in the system, we compute the index $\langle\Delta_{3}(L)\rangle$, defined as the average of $\Delta_{3}(L)$ over the range $L\in[10,40]$. For reference, the typical values of $\langle\Delta_{3}\rangle$ in the three limiting cases are approximately $\langle\Delta_{3}\rangle_{\rm Poisson}\simeq 1.67$, $\langle\Delta_{3}\rangle_{\rm GOE}\simeq 0.319$, and $\langle\Delta_{3}\rangle_{\rm GUE}\simeq 0.184$.
For comparison, we fit the level spacing distribution between Poisson and GOE taking the semi-Poisson form  [bogomolny2009spectral, bogomolny2001short, bogomolny2004spectral]
$P_{\rm SP}(\beta,s)=A_{\beta}\,s^{\beta}\,\exp\!\left[-(\beta+1)s\right],\qquad A_{\beta}=\frac{(\beta+1)^{\beta+1}}{\Gamma(\beta+1)}.$ (5)
We also fit the level spacing distribution between GOE and GUE taking the form [mehta2004random]
$P_{\rm WD}(\gamma,s)=a_{\gamma}\,s^{\gamma}\,\exp\!\left(-b_{\gamma}s^{2}\right),$ (6)
where we have defined
$b_{\gamma}=\left[\frac{\Gamma\!\left(\frac{\gamma+2}{2}\right)}{\Gamma\!\left(\frac{\gamma+1}{2}\right)}\right]^{2}\quad\textrm{and}\quad a_{\gamma}=\frac{2\,b_{\gamma}^{(\gamma+1)/2}}{\Gamma\!\left(\frac{\gamma+1}{2}\right)}.$ (7)
with $\gamma=1$ and $\gamma=2$ correspond to GOE and GUE, respectively.
VI.5 Correlation function analysis
To quantify how rotation influences the overall structure of the system’s eigenstates, we analyze the two-point correlation function,
$C(\mathbf{r},\mathbf{r}^{\prime})=\langle\psi^{*}(\mathbf{r})\psi(\mathbf{r}^{\prime})\rangle.$ (8)
According to Berry’s random wave model for chaotic cavities, the wave function can be modeled as a superposition of plane waves with random phases and a fixed wavenumber $k$, $\psi(\mathbf{r})=\sum_{n=1}^{N}a_{n}e^{i\mathbf{k}_{n}\cdot\mathbf{r}}$. After ensemble averaging, the correlation function becomes isotropic and depends only on the separation $r=|\mathbf{r}-\mathbf{r}^{\prime}|$, taking the universal form
$C(r)=J_{0}(kr),$ (9)
where $J_{0}$ is the zeroth-order Bessel function of the first kind [berry1977regular].
In contrast, in our model the correlation function is generally anisotropic and depends on both the magnitude and direction of the separation vector. To extract a scalar measure from the anisotropic correlation function, we define
$C(r)\equiv\big\langle\max_{\hat{\mathbf{r}}^{\prime}}C(\bm{\rho},\bm{\rho}+\mathbf{r}^{\prime})\big\rangle_{\bm{\rho}},\qquad|\mathbf{r}^{\prime}|=r,$ (10)
where the maximization is performed over all directions $\hat{\mathbf{r}}^{\prime}=\mathbf{r}^{\prime}/|\mathbf{r}^{\prime}|$ of the displacement vector $\mathbf{r}^{\prime}$, for a fixed separation $|\mathbf{r}^{\prime}|=r$, and $\langle\cdots\rangle_{\bm{\rho}}$ denotes an average over initial positions $\bm{\rho}$. We analyze the resulting correlation function by fitting it to the form
$C(r)=J_{0}\!\left(2\pi\frac{r}{l}\right),$ (11)
where $l$ defines a correlation length that characterizes the spatial correlations of the eigenstates. We emphasize that the correlation length is used here as a relative and heuristic measure of spatial correlations. For a fully chaotic system, the correlation length reduces to the wavelength $\lambda$ of the eigenfunction.
VI.6 Semiclassical ray-dynamics
Direct time-stepping ray propagation in polygonal billiards with anisotropic dispersion is prone to numerical issues near grazing incidence and near vertices, where corner sensitivity can lead to trajectory-dependent ambiguities. To obtain stable long-time dynamics, we implement the semiclassical motion as a discrete map between successive boundary collisions. We tabulate the relation between the incidence angle $\varphi_{\mathrm{in}}$ and the outgoing angle $\varphi_{\mathrm{out}}$, determined by the conservation of energy and the tangential component $k_{t}$. Each table contains about $3\times 10^{5}$ sampled angles. This approach encodes the warped reflection law once and for all, and removes the need to solve for $\varphi_{\mathrm{out}}$ by conducting a local root search at every collision.
Between collisions, the next collision point is found by ray-boundary intersection. The dynamics is thus an alternation of (i) intersection updates for free flight and (ii) table-based updates for reflection. This map-based scheme maintains high angular accuracy even for nearly grazing trajectories and avoids artificial corner sticking and other corner-related numerical artifacts. As a result, we can compute long trajectories and Poincaré sections robustly.
Acknowledgements.
This project was supported by the National Science Foundation (Grant No. 2403491). We are thankful for the useful discussions with E. Räsänen, A. M. Bozkur, and J. Velasco Jr. A.M.G. thanks the Studienstiftung des Deutschen Volkes for financial support.